A new roller coaster at an amusement park requires individuals to be at least​ 4' 8" ​(56 inches) tall to ride. It is estimated that the heights of​ 10-year-old boys are normally distributed with μ=54.0 inches and σ=5 inches.

a. What proportion of​ 10-year-old boys is tall enough to ride the​ coaster?

b. A smaller coaster has a height requirement of 50 inches to ride. What proportion of​ 10-year-old boys is tall enough to ride this​ coaster?

c. What proportion of​ 10-year-old boys is tall enough to ride the coaster in part b but not tall enough to ride the coaster in part​ a?

a. The proportion of​ 10-year-old boys tall enough to ride the coaster is

nothing. ​(Round to four decimal places as​ needed.)

b. The proportion of 10​ year-old-boys tall enough to ride the smaller coaster is

nothing. (Round to four decimal places as​ needed.)

c. The proportion of​ 10-year-old boys tall enough to ride the coaster in part b but not tall enough to ride the coaster in part a is nothing.
Answer #1

Mean = 54 inches S.D = 5

a)

X = 56

We need to find the z-score here.

o = Z n - X

z = (56 - 54) / 5 = 2/5 = 0.4

Which gives a value of 0.6554 in the z-table.

This proportion is for the kids less than 56 inches. For more than 56 inches, subtract it with 1.

= 1 - 0.6554 = 0.3446

The answer is 0.3446.

b)

When X = 50

z = (50 - 54) / 5 = -4/5 = -0.8

Which gives a value of 0.2119 in the z-table.

This proportion is for the kids less than 50 inches. For more than 50 inches, subtract it with 1.

= 1 - 0.2119 = 0.7881

The answer is 0.7881.

c)

For this, we need to subtract the proportion we got in part b with part a.

= 0.7881 - 0.3446 = 0.4435

The answer is 0.4435.